OpenStudy (anonymous):

105^2-8^2= Factor this for fan and medal.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

\[a^2-b^2=(a-b)(a+b) \text{is a difference of squares factored }\]

OpenStudy (anonymous):

so im guessing it would be 105^2-8^2=(105-8)(105+8)